# Crosscap states and duality of Ising field theory in two dimensions

### v1

@@ Abstract
$\prod_{\alpha=1}^{M}b_{-n_{\alpha}+1/2}\prod_{\alpha=1}^{M}\bar{b}_{-n_{\alpha}+1/2}=\begin{cases}(-1)^{\frac{M}{2}}\prod_{\alpha=1}^{M}b_{-n_{\alpha}+1/2}\bar{b}_{-n_{\alpha}+1/2}&\quad M\;\mathrm{even}\\ (-1)^{\frac{M-1}{2}}\prod_{\alpha=1}^{M}b_{-n_{\alpha}+1/2}\bar{b}_{-n_{\alpha}+1/2}&\quad M\;\mathrm{odd}\\ \end{cases}\,.$ (S71)
Therefore, we find that $|\mathcal{C}_{\pm}\rangle$ are the following linear combinations of the crosscap Ishibashi states $|1\rangle\rangle_{\mathcal{C}}$ and $|\varepsilon\rangle\rangle_{\mathcal{C}}$:
$|\mathcal{C}_{\pm}\rangle=\sqrt{\frac{2+\sqrt{2}}{2}}|1\rangle\rangle_{\mathcal{C}}\pm\sqrt{\frac{2-\sqrt{2}}{2}}|\varepsilon\rangle\rangle_{\mathcal{C}}\,,$ (S72)
both of which are physical conformal crosscap states satisfying the loop channel-tree channel equivalence between the Klein bottle partition function and the crosscap partition function $\langle\mathcal{C}_{\pm}|e^{-\beta H_{0}}|\mathcal{C}_{\pm}\rangle$ [22]. This corresponds to Eq.(11) in the main text.
Lastly, we note that the notation used here differs slightly from that in Ref. [40]. Specifically, we identify $\sigma_{j}^{x}$ as the lattice realization of the spin field $\sigma$ and $-\sigma_{j}^{z}$ as the lattice realization of the energy field $\varepsilon$, with the fermionic representation $\varepsilon=-i\chi\bar{\chi}$. This choice ensures that the operator product expansion (OPE) coefficient $C_{\sigma\sigma}^{\varepsilon}$ is positive. Additionally, the Majorana fields $\chi(x)$ and $\bar{\chi}(x)$ in Eq. (S55) correspond to the lattice realizations $\chi_{j}$ and $\bar{\chi}_{j}$ in Eq. (S9), respectively.
D.3 C. Exact crosscap overlaps in the presence of thermal perturbation
@@
As a side remark, we note that the above expression of $\Theta(s_{1})$ is expanded in orders of the energy spectrum, which indicates that the truncated conformal space approach may serve as an effective non-perturbative method for calculating the crosscap overlap. In fact, by retaining about one hundred terms in the expansion of $\Theta(s_{1})$, one can already obtain the universal scaling function for the crosscap overlap with very high precision.
@@
Viewed as the continuous deformation from the (unperturbed) CFT eigenstates $|\psi_{n_{1}\cdots n_{M}}(0)\rangle=\prod_{\alpha=1}^{M}b_{-n+1/2}\bar{b}_{-n+1/2}|0\rangle_{\mathrm{NS}}$, the perturbed eigenstates
@@
where the calculation is similar to the lattice case (see Sec. C.3). This is the continuum counterpart of the lattice crosscap overlaps in Eq. (S50). In the off-critical region, where $h\neq 1$ ($s_{1}\neq 0$), finite-size corrections are present.
With the above crosscap overlap, one can expand the crosscap state $|\mathcal{C}_{\pm}\rangle$ in the eigenbasis of the Ising CFT with the thermal perturbation:
$|\mathcal{C}_{\pm}\rangle=\frac{e^{i\left[\frac{\pi}{8}\pm\Theta(s_{1})\right]}}{\sqrt{2}}\exp\left[\pm i\sum_{n=1}^{\infty}(-1)^{n}\eta^{\dagger}_{n-\frac{1}{2}}\eta^{\dagger}_{-n+\frac{1}{2}}\right]|\psi_{0}(s_{1})\rangle+\frac{e^{-i\left[\frac{\pi}{8}\pm\Theta(s_{1})\right]}}{\sqrt{2}}\exp\left[\mp i\sum_{n=1}^{\infty}(-1)^{n}\eta^{\dagger}_{n-\frac{1}{2}}\eta^{\dagger}_{-n+\frac{1}{2}}\right]|\psi_{0}(s_{1})\rangle\,.$ (S86)
@@
$\bar{\Psi}(\bar{z})$ $:\displaystyle=\sqrt{\frac{2\pi\bar{w}}{L}}:e^{i\bar{\phi}(\bar{w})}:\,,\quad\bar{\Psi}^{\dagger}(\bar{z})=\sqrt{\frac{2\pi\bar{w}}{L}}:e^{-i\bar{\phi}(\bar{w})}$ : (S88)
@@
The momentum modes satisfy commutation relations $[x_{0},a_{0}]=[\bar{x}_{0},\bar{a}_{0}]=i$ and $[a_{k},a_{l}]=[\bar{a}_{k},\bar{a}_{l}]=k\delta_{k+l}\;\forall k,l\in\mathbb{Z}$. The Virasoro primary states of the compactified boson CFT are labeled by the eigenvalues of $a_{0}$ and $\bar{a}_{0}$:
@@
This is equivalent to the condition $[\chi(x)\mp i\bar{\chi}(x+L/2)]|B_{\pm}\rangle=0$, where $\chi$ and $\bar{\chi}$ are the Majorana fields [Eq. (S55)] in the Ising CFT. Consequently, for two copies of such fermionic Gaussian states $|B_{\pm}\rangle$, we obtain the constraints on the Dirac fields:
@@
and the zero mode part is fixed as
@@
with the normalization $\langle 0,0|(|B_{\pm}\rangle^{(1)}|B_{\mp}\rangle^{(2)})=({}_{\mathrm{NS}}\langle 0|B_{+}\rangle)({}_{\mathrm{NS}}\langle 0|B_{-}\rangle)=1$.
@@
The naive two copies of Ising crosscap states $|\mathcal{C}_{\pm}\rangle^{(1)}|\mathcal{C}_{\pm}\rangle^{(2)}$are not invariant under the action of $G$. However, the two copies of the Ising Ishibashi states $|\alpha\rangle\rangle_{\mathcal{C}}^{(1)}|\alpha\rangle\rangle_{\mathcal{C}}^{(2)}$, with $\alpha=1,\varepsilon$, are suitable crosscap states for the $\mathbb{Z}_{2}$ orbifold boson CFT:
$|1\rangle\rangle_{\mathcal{C}}^{(1)}|1\rangle\rangle_{\mathcal{C}}^{(2)}$ $=\left[\frac{1}{2}(|B_{+}\rangle^{(1)}+|B_{-}\rangle^{(1)})\right]\left[\frac{1}{2}(|B_{+}\rangle^{(2)}+|B_{-}\rangle^{(2)})\right]=\frac{1}{2}\left(|\mathcal{O}_{+}\rangle+|\mathcal{O}_{-}\rangle\right)\,,$
@@
The bosonization dictionary for two copies of $\varepsilon$ fields and $\sigma$ fields is given by [12]:
@@
where $\langle\partial\phi(w)\partial\phi(w^{\prime})\rangle=-\frac{1}{(w-w^{\prime})^{2}}$ and ¯$\langle\bar{\partial}\bar{\phi}(\bar{w})\bar{\partial}\bar{\phi}(\bar{w}^{\prime})\rangle=-\frac{1}{(\bar{w}-\bar{w}^{\prime})^{2}}$ are the (anti-)chiral current-current correlators on the plane, respectively. For the vertex operators, the theorem gives
@@
where we used the commutation relations $[\phi^{\prime}_{-}(w),a_{-k}]=iw^{-k}$ and $[\bar{\phi}^{\prime}_{-}(\bar{w}),a_{-k}]=i\bar{w}^{-k}$, $\forall k>0$.
Theorem 1 [Eq. (S112)] is a corollary of Lemma 1. Since $[\partial\phi^{\prime}_{-}(w),K]=-\frac{1}{w^{2}}\bar{\partial}\bar{\phi}^{\prime}_{+}(-1/w)$ and [¯$[\bar{\partial}\bar{\phi}^{\prime}_{-}(\bar{w}),K]=-\frac{1}{\bar{w}^{2}}\partial\phi^{\prime}_{+}(-1/\bar{w})$, we have
@@
where in the second and fourth equalities, we used the Baker–Campbell–Hausdorff formula: $e^{A}e^{B}=e^{A+B+\frac{1}{2}[A,B]}=e^{B}e^{A}e^{[A,B]}$, assuming $[[A,B],A]=[[A,B],B]=0$; in the third and fifth equalities, we used the Lemma 1 [Eq. (S116)]; and in the last equality, we used the correlator $[\phi^{\prime}_{-}(w),\phi^{\prime}_{+}(-1/\bar{w})]=\langle\phi^{\prime}(w)\phi^{\prime}(-1/\bar{w})\rangle=-\ln(1+\frac{1}{w\bar{w}})$.
@@
×$\quad\times e^{i\alpha_{2}\vartheta^{\prime}_{-}(w_{2},\bar{w}_{2})}e^{\pm i\alpha_{1}\vartheta^{\prime}_{+}(-1/\bar{w}_{1},-1/w_{1})}e^{i\alpha_{1}\vartheta^{\prime}_{-}(w_{1},\bar{w}_{1})}$
@@
×$\quad\times\left(e^{\pm i\alpha_{1}\vartheta^{\prime}_{+}(-1/\bar{w}_{1},-1/w_{1})}e^{i\alpha_{2}\vartheta^{\prime}_{-}(w_{2},\bar{w}_{2})}e^{\mp\alpha_{1}\alpha_{2}\langle\vartheta^{\prime}(w_{2},\bar{w}_{2})\vartheta^{\prime}(-1/\bar{w}_{1},-1/w_{1})\rangle}\right)e^{i\alpha\vartheta^{\prime}_{-}(w_{1},\bar{w}_{1})}$
@@
×$\quad\times e^{\pm K\pm i\alpha_{1}\vartheta^{\prime}_{+}(-1/\bar{w}_{1},-1/w_{1})\pm i\alpha_{2}\vartheta^{\prime}_{+}(-1/\bar{w}_{2},-1/w_{2})}:e^{i\alpha_{1}\vartheta^{\prime}(w_{1},\bar{w}_{1})+i\alpha_{2}\vartheta^{\prime}(w_{2},\bar{w}_{2})}:$
@@
where, in the second equality, we used the result for the $n=1$ case [Eq. (S121)] by induction. We also used the definition of the normal-ordered vertex operators [Eq. (S93)] and the Baker–Campbell–Hausdorff formula [Below Eq. (S120)], along with the correlator
@@
${}_{\mathrm{NS}}\langle 0|\prod_{j=1}^{n}\varepsilon(w_{j},\bar{w}_{j})|\mathcal{C}_{\pm}\rangle=\sqrt{\frac{2+\sqrt{2}}{2}}\;{}_{\mathrm{NS}}\langle 0|\prod_{j=1}^{n}\varepsilon(w_{j},\bar{w}_{j})|1\rangle\rangle_{\mathcal{C}}\pm\sqrt{\frac{2-\sqrt{2}}{2}}\;{}_{\mathrm{NS}}\langle 0|\prod_{j=1}^{n}\varepsilon(w_{j},\bar{w}_{j})|\varepsilon\rangle\rangle_{\mathcal{C}}\,,$ (S124)
where the bosonization of the crosscap correlator in each Ishibashi sector is given by:
${}_{\mathrm{NS}}\langle 0|\prod_{j=1}^{n}\varepsilon(w_{j},\bar{w}_{j})|1\rangle\rangle_{\mathcal{C}}$ $=\sqrt{\frac{1}{2}\langle 0,0|\prod_{j=1}^{n}\partial\phi(w_{j})\bar{\partial}\bar{\phi}(\bar{w}_{j})\left(|\mathcal{O}_{+}\rangle+|\mathcal{O}_{-}\rangle\right)}\,,$
@@
×$\quad\times\langle 0,0|:\prod_{j\neq k_{1},\ldots,k_{2p}}\partial\phi(w_{j})\prod_{l\neq k^{\prime}_{1},\ldots,k^{\prime}_{2q}}\bar{\partial}\bar{\phi}(\bar{w}_{l}):e^{\pm K}|0,0\rangle\,,$ (S127)
@@
and $[\frac{n}{2}]$ is the floor function. Next, using the generalized Wick’s theorem Eq. (S112), the expression simplifies further:
@@
=(±$=(\pm)^{n}\sum_{0\leq p\leq\left[\frac{n}{2}\right]}\sum_{\begin{subarray}{c}1\leq k_{1}<\cdots<k_{2p}\leq n\\ 1\leq k^{\prime}_{1}<\cdots<k^{\prime}_{2p}\leq n\end{subarray}}\mathrm{Hf}\left[\frac{1}{(w_{k_{\alpha}}-w_{k_{\beta}})^{2}}\right]\mathrm{Hf}\left[\frac{1}{(w_{k^{\prime}_{\alpha}}-w_{k^{\prime}_{\beta}})^{2}}\right]_{\alpha,\beta=1,\ldots,2p}\mathrm{Perm}\left[\frac{1}{(1+w_{j}\bar{w}_{l})^{2}}\right]_{\begin{subarray}{c}j\neq k_{1},\ldots,k_{2p}\\ l\neq k^{\prime}_{1},\ldots,k^{\prime}_{2p}\end{subarray}}\,,$ (S134)
@@
×$\times\sqrt{\sum_{0\leq p\leq\left[\frac{n}{2}\right]}\sum_{\begin{subarray}{c}1\leq k_{1}<\cdots<k_{2p}\leq n\\ 1\leq k^{\prime}_{1}<\cdots<k^{\prime}_{2p}\leq n\end{subarray}}\mathrm{Hf}\left[\frac{1}{(w_{k_{\alpha}}-w_{k_{\beta}})^{2}}\right]\mathrm{Hf}\left[\frac{1}{(w_{k^{\prime}_{\alpha}}-w_{k^{\prime}_{\beta}})^{2}}\right]_{\alpha,\beta=1,\ldots,2p}\mathrm{Perm}\left[\frac{1}{(1+w_{j}\bar{w}_{l})^{2}}\right]_{\begin{subarray}{c}j\neq k_{1},\ldots,k_{2p}\\ l\neq k^{\prime}_{1},\ldots,k^{\prime}_{2p}\end{subarray}}}$ (S135)
In the end, we present the results for the one- and two-point correlators as examples:
For the one-point correlator, we have:
${}_{\mathrm{NS}}\langle 0|\varepsilon(w,\bar{w})|\mathcal{C}_{\pm}\rangle$ =±$=\pm\sqrt{\frac{2-\sqrt{2}}{2}}\frac{1}{1+w\bar{w}}\,,$ (S136)
@@
According to the fusion rule $\sigma\sigma\sim 1+\varepsilon$, $(2n+1)$-point crosscap correlators of the $\sigma$ field vanish. For $2n$-point crosscap correlators, the calculation follows a similar logic to that of the $\varepsilon$ field. We express the full crosscap correlator as a combination of two Ishibashi crosscap correlators:
${}_{\mathrm{NS}}\langle 0|\prod_{j=1}^{2n}\sigma(w_{j},\bar{w}_{j})|\mathcal{C}_{\pm}\rangle=\sqrt{\frac{2+\sqrt{2}}{2}}{}_{\mathrm{NS}}\langle 0|\prod_{j=1}^{2n}\sigma(w_{j},\bar{w}_{j})|1\rangle\rangle_{\mathcal{C}}\pm\sqrt{\frac{2-\sqrt{2}}{2}}{}_{\mathrm{NS}}\langle 0|\prod_{j=1}^{2n}\sigma(w_{j},\bar{w}_{j})|\varepsilon\rangle\rangle_{\mathcal{C}}\,.$ (S139)
@@
${}_{\mathrm{NS}}\langle 0|\prod_{j=1}^{2n}\sigma(w_{j},\bar{w}_{j})|1\rangle\rangle_{\mathcal{C}}$ $=\sqrt{2^{n-1}\langle 0,0|\prod_{j=1}^{2n}\cos\frac{\vartheta}{2}(w_{j},\bar{w}_{j})\left(|\mathcal{O}_{+}\rangle+|\mathcal{O}_{-}\rangle\right)}\,,$
@@
$=\langle 0,-\frac{1}{2}\sum_{j=1}^{2n}\epsilon_{j}|\mathcal{O}^{0}_{\pm}\rangle\cdot\exp\left[-\frac{1}{4}\sum_{j=1}^{2n}\epsilon_{j}\cdot\sum_{l=1}^{2n}\epsilon_{l}\ln(w_{l}\bar{w}_{l})\right]\,,$ (S145)
@@
×$\times\prod_{j=1}^{2n}|w_{j}|^{-2m\epsilon_{j}}\prod_{1\leq j<l\leq 2n}|w_{j}-w_{l}|^{\epsilon_{j}\epsilon_{l}/2}\prod_{1\leq j,l\leq 2n}\left(1+\frac{1}{w_{j}\bar{w}_{l}}\right)^{-\epsilon_{j}\epsilon_{l}/4}$

### v2

@@ Abstract
$\prod_{\alpha=1}^{M}b_{-n_{\alpha}+1/2}\prod_{\alpha=1}^{M}\bar{b}_{-n_{\alpha}+1/2}=\begin{cases}(-1)^{\frac{M}{2}}\prod_{\alpha=1}^{M}b_{-n_{\alpha}+1/2}\bar{b}_{-n_{\alpha}+1/2}&\quad M\;\mathrm{even}\\ (-1)^{\frac{M-1}{2}}\prod_{\alpha=1}^{M}b_{-n_{\alpha}+1/2}\bar{b}_{-n_{\alpha}+1/2}&\quad M\;\mathrm{odd}\end{cases}.$ (S89)
Therefore, $|\mathcal{C}_{\pm}\rangle$ are the following linear combinations of the crosscap Ishibashi states:
$|\mathcal{C}_{\pm}\rangle=\sqrt{\frac{2+\sqrt{2}}{2}}|\mathbbm{1}\rangle\rangle_{\mathcal{C}}\pm\sqrt{\frac{2-\sqrt{2}}{2}}|\varepsilon\rangle\rangle_{\mathcal{C}}\,.$ (S90)
This corresponds to Eq. (13) in the main text. Comparing with Eqs. (S75) and (S76), we find that $|\mathcal{C}_{\pm}\rangle$ are the (standard) PSS conformal crosscap state $|\mathcal{C}_{\mathbbm{1}}\rangle$ and its dual $\frac{\mathcal{D}_{\sigma}}{\langle\mathcal{D}_{\sigma}\rangle}|\mathcal{C}_{\mathbbm{1}}\rangle$, respectively.
Moreover, from Eq. (S67), we can derive the continuum counterpart of the modified lattice crosscap states $|\mathcal{C}^{\prime}_{\mathrm{latt}}\rangle$:
$|\mathcal{C}^{\prime}\rangle$ $=\frac{-ie^{i\pi/8}}{\sqrt{2}}|B_{+}\rangle+\frac{ie^{-i\pi/8}}{\sqrt{2}}|B_{-}\rangle$
$=\sqrt{\frac{2-\sqrt{2}}{2}}|\mathbbm{1}\rangle\rangle_{\mathcal{C}}-\sqrt{\frac{2+\sqrt{2}}{2}}|\varepsilon\rangle\rangle_{\mathcal{C}}\,$, (S91)
where we used Eqs. (S87) and (S88). This proves $|\mathcal{C}^{\prime}\rangle=|\mathcal{C}_{\varepsilon}\rangle$, where $|\mathcal{C}_{\varepsilon}\rangle$ is one of the PSS crosscap states in Eq. (S75).
Having identified the continuum counterparts of $|\mathcal{C}_{\mathrm{latt}}^{\pm}\rangle$ and $|\mathcal{C}^{\prime}_{\mathrm{latt}}\rangle$ in the Ising CFT, we find that Eq. (S34) is the perfect lattice correspondence of Eq. (S76), which can be interpreted as the interplay of the Kramers-Wannier TDL and the PSS crosscap states.
D.4 D. Exact crosscap overlaps in the presence of thermal perturbation
@@
As a side remark, we note that the above expression of $\Theta(s_{1})$ is expanded in orders of the energy levels. This indicates that the truncated conformal space approach may serve as an effective non-perturbative method for calculating the crosscap overlap. In fact, by retaining about one hundred terms in the expansion of $\Theta(s_{1})$, one can already obtain the universal scaling function for the crosscap overlap with very high precision.
@@
Viewed as a continuous deformation of the (unperturbed) CFT eigenstates $|\psi_{n_{1}\cdots n_{M}}(0)\rangle=\prod_{\alpha=1}^{M}b_{-n+1/2}\bar{b}_{-n+1/2}|0\rangle_{\mathrm{NS}}$, the perturbed eigenstates
@@
where the calculation is similar to the lattice case (see Sec. C.3). This is the continuum counterpart of the lattice crosscap overlaps in Eq. (S58). However, we note that in the off-critical region of the lattice model ($h\neq 1$), finite-size corrections are present, unlike the critical point ($h=1$) where finite-size corrections are absent.
In summary, we have the expansion of the crosscap state $|\mathcal{C}_{\pm}\rangle$ in the eigenbasis of the Ising CFT with the thermal perturbation:
$|\mathcal{C}_{\pm}\rangle=\frac{e^{i\left[\frac{\pi}{8}\pm\Theta(s_{1})\right]}}{\sqrt{2}}\exp\left[\pm i\sum_{n=1}^{\infty}(-1)^{n}\eta^{\dagger}_{n-\frac{1}{2}}\eta^{\dagger}_{-n+\frac{1}{2}}\right]|\psi_{0}(s_{1})\rangle+\frac{e^{-i\left[\frac{\pi}{8}\pm\Theta(s_{1})\right]}}{\sqrt{2}}\exp\left[\mp i\sum_{n=1}^{\infty}(-1)^{n}\eta^{\dagger}_{n-\frac{1}{2}}\eta^{\dagger}_{-n+\frac{1}{2}}\right]|\psi_{0}(s_{1})\rangle\,$, (S105)
where $\Theta(s_{1})$ has been defined in Eq. (S101). More specifically, the squared norms of the crosscap overlaps take the simple form,
$|\langle\psi_{n_{1}\cdots n_{M}}(s_{1})|\mathcal{C}_{\pm}\rangle|^{2}=1+\frac{(-1)^{M}}{\sqrt{1+e^{\mp 2\pi s_{1}}}}\,$, (S106)
where $|\psi_{n_{1}\cdots n_{M}}(s_{1})\rangle$ is the eigenstate deformed from the unperturbed state $|\psi_{n_{1}\cdots n_{M}}(0)\rangle\equiv\prod_{\alpha=1}^{M}b_{-n_{\alpha}+1/2}\bar{b}_{-n_{\alpha}+1/2}|0\rangle_{\mathrm{NS}}$ in the Ising CFT basis.
@@
$\bar{\Psi}(\bar{z})$ $=\sqrt{\frac{2\pi\bar{w}}{L}}:e^{i\bar{\phi}(\bar{w})}:\,,\quad\bar{\Psi}^{\dagger}(\bar{z})=\sqrt{\frac{2\pi\bar{w}}{L}}:e^{-i\bar{\phi}(\bar{w})}:$ (S108)
@@
The momentum modes satisfy commutation relations $x_{0},a_{0}]=[\bar{x}_{0},\bar{a}_{0}]=i$ and $a_{k},a_{l}]=[\bar{a}_{k},\bar{a}_{l}]=k\delta_{k+l}\;\forall k,l\in\mathbb{Z}$. The Virasoro primary states of the compactified boson CFT are labeled by the eigenvalues of $a_{0}$ and $\bar{a}_{0}$:
@@
This is equivalent to the condition $\chi(x)\mp i\bar{\chi}(x+L/2)]|B_{\pm}\rangle=0$, where $\chi$ and $\bar{\chi}$ are the Majorana fields [Eq. (S78)] in the Ising CFT. Consequently, for two copies of such fermionic Gaussian states $|B_{\pm}\rangle$, we obtain the constraints on the Dirac fields:
@@
and the zero-mode part is fixed as
@@
with the normalization $\langle 0,0|(|B_{\pm}\rangle^{(1)}|B_{\mp}\rangle^{(2)})={}_{\mathrm{NS}}\langle 0|B_{+}\rangle\cdot{}_{\mathrm{NS}}\langle 0|B_{-}\rangle=1$.
@@
We find that the two copies of the Ising Ishibashi states $|\alpha\rangle\rangle_{\mathcal{C}}^{(1)}|\alpha\rangle\rangle_{\mathcal{C}}^{(2)}$, with $\alpha=\mathbbm{1},\varepsilon$, are suitable crosscap states for the $\mathbb{Z}_{2}$ orbifold boson CFT:
$|\mathbbm{1}\rangle\rangle_{\mathcal{C}}^{(1)}|\mathbbm{1}\rangle\rangle_{\mathcal{C}}^{(2)}$ $=\left[\frac{1}{2}(|B_{+}\rangle^{(1)}+|B_{-}\rangle^{(1)})\right]\left[\frac{1}{2}(|B_{+}\rangle^{(2)}+|B_{-}\rangle^{(2)})\right]=\frac{1}{2}\left(|\mathcal{O}_{+}\rangle+|\mathcal{O}_{-}\rangle\right)\,$,
@@
The bosonization dictionary for two copies of $\varepsilon$ fields and $\sigma$ fields is given by [12]
@@
where $\langle\partial\phi(w)\partial\phi(w^{\prime})\rangle=-\frac{1}{(w-w^{\prime})^{2}}$ and $\langle\bar{\partial}\bar{\phi}(\bar{w})\bar{\partial}\bar{\phi}(\bar{w}^{\prime})\rangle=-\frac{1}{(\bar{w}-\bar{w}^{\prime})^{2}}$ are the (anti-)chiral current-current correlators on the plane, respectively. For the vertex operators, the theorem gives
@@
where we used the commutation relations $\phi^{\prime}_{-}(w),a_{-k}]=iw^{-k}$ and $\bar{\phi}^{\prime}_{-}(\bar{w}),a_{-k}]=i\bar{w}^{-k}$, $\forall k>0$.
Theorem 1 [Eq. (S132)] is a corollary of Lemma 1. Since $\partial\phi^{\prime}_{-}(w),K]=-\frac{1}{w^{2}}\bar{\partial}\bar{\phi}^{\prime}_{+}(-1/w)$ and $\bar{\partial}\bar{\phi}^{\prime}_{-}(\bar{w}),K]=-\frac{1}{\bar{w}^{2}}\partial\phi^{\prime}_{+}(-1/\bar{w})$, we have
@@
where in the second and fourth equalities, we used the Baker-Campbell-Hausdorff formula: $e^{A}e^{B}=e^{A+B+\frac{1}{2}[A,B]}=e^{B}e^{A}e^{[A,B]}$, assuming $[[A,B],A]=[[A,B],B]=0$; in the third and fifth equalities, we used the Lemma 1 [Eq. (S136)]; and in the last equality, we used the correlator $\phi^{\prime}_{-}(w),\phi^{\prime}_{+}(-1/\bar{w})]=\langle\phi^{\prime}(w)\phi^{\prime}(-1/\bar{w})\rangle=-\ln(1+\frac{1}{w\bar{w}})$.
@@
$\quad\times e^{i\alpha_{2}\vartheta^{\prime}_{-}(w_{2},\bar{w}_{2})}e^{\pm i\alpha_{1}\vartheta^{\prime}_{+}(-1/\bar{w}_{1},-1/w_{1})}e^{i\alpha_{1}\vartheta^{\prime}_{-}(w_{1},\bar{w}_{1})}$
@@
$\quad\times\left(e^{\pm i\alpha_{1}\vartheta^{\prime}_{+}(-1/\bar{w}_{1},-1/w_{1})}e^{i\alpha_{2}\vartheta^{\prime}_{-}(w_{2},\bar{w}_{2})}e^{\mp\alpha_{1}\alpha_{2}\langle\vartheta^{\prime}(w_{2},\bar{w}_{2})\vartheta^{\prime}(-1/\bar{w}_{1},-1/w_{1})\rangle}\right)e^{i\alpha\vartheta^{\prime}_{-}(w_{1},\bar{w}_{1})}$
@@
$\quad\times e^{\pm K\pm i\alpha_{1}\vartheta^{\prime}_{+}(-1/\bar{w}_{1},-1/w_{1})\pm i\alpha_{2}\vartheta^{\prime}_{+}(-1/\bar{w}_{2},-1/w_{2})}:e^{i\alpha_{1}\vartheta^{\prime}(w_{1},\bar{w}_{1})+i\alpha_{2}\vartheta^{\prime}(w_{2},\bar{w}_{2})}:$
@@
where, in the second equality, we used the result for the $n=1$ case [Eq. (S141)] by induction. We also used the definition of the normal-ordered vertex operators [Eq. (S113)] and the Baker-Campbell-Hausdorff formula [Below Eq. (S140)], along with the correlator
@@
${}_{\mathrm{NS}}\langle 0|\prod_{j=1}^{n}\varepsilon(w_{j},\bar{w}_{j})|\mathcal{C}_{\pm}\rangle=\sqrt{\frac{2+\sqrt{2}}{2}}\;{}_{\mathrm{NS}}\langle 0|\prod_{j=1}^{n}\varepsilon(w_{j},\bar{w}_{j})|\mathbbm{1}\rangle\rangle_{\mathcal{C}}\pm\sqrt{\frac{2-\sqrt{2}}{2}}\;{}_{\mathrm{NS}}\langle 0|\prod_{j=1}^{n}\varepsilon(w_{j},\bar{w}_{j})|\varepsilon\rangle\rangle_{\mathcal{C}}\,$, (S144)
where the bosonization of the crosscap correlator in each Ishibashi sector is given by
${}_{\mathrm{NS}}\langle 0|\prod_{j=1}^{n}\varepsilon(w_{j},\bar{w}_{j})|\mathbbm{1}\rangle\rangle_{\mathcal{C}}$ $=\sqrt{\frac{1}{2}\langle 0,0|\prod_{j=1}^{n}\partial\phi(w_{j})\bar{\partial}\bar{\phi}(\bar{w}_{j})\left(|\mathcal{O}_{+}\rangle+|\mathcal{O}_{-}\rangle\right)}\,$,
@@
$\quad\times\langle 0,0|:\prod_{j\neq k_{1},\ldots,k_{2p}}\partial\phi(w_{j})\prod_{l\neq k^{\prime}_{1},\ldots,k^{\prime}_{2q}}\bar{\partial}\bar{\phi}(\bar{w}_{l}):e^{\pm K}|0,0\rangle\,$, (S147)
@@
and $\frac{n}{2}]$ denotes the integer part of $\frac{n}{2}$. Next, using the generalized Wick’s theorem in Eq. (S132), the expression simplifies further:
@@
$=(\pm)^{n}\sum_{0\leq p\leq\left[\frac{n}{2}\right]}\sum_{\begin{subarray}{c}1\leq k_{1}<\cdots<k_{2p}\leq n\\ 1\leq k^{\prime}_{1}<\cdots<k^{\prime}_{2p}\leq n\end{subarray}}\mathrm{Hf}\left[\frac{1}{(w_{k_{\alpha}}-w_{k_{\beta}})^{2}}\right]\mathrm{Hf}\left[\frac{1}{(w_{k^{\prime}_{\alpha}}-w_{k^{\prime}_{\beta}})^{2}}\right]_{\alpha,\beta=1,\ldots,2p}\mathrm{Perm}\left[\frac{1}{(1+w_{j}\bar{w}_{l})^{2}}\right]_{\begin{subarray}{c}j\neq k_{1},\ldots,k_{2p}\\ l\neq k^{\prime}_{1},\ldots,k^{\prime}_{2p}\end{subarray}}\,.$ (S154)
@@
$\times\sqrt{\sum_{0\leq p\leq\left[\frac{n}{2}\right]}\sum_{\begin{subarray}{c}1\leq k_{1}<\cdots<k_{2p}\leq n\\ 1\leq k^{\prime}_{1}<\cdots<k^{\prime}_{2p}\leq n\end{subarray}}\mathrm{Hf}\left[\frac{1}{(w_{k_{\alpha}}-w_{k_{\beta}})^{2}}\right]\mathrm{Hf}\left[\frac{1}{(w_{k^{\prime}_{\alpha}}-w_{k^{\prime}_{\beta}})^{2}}\right]_{\alpha,\beta=1,\ldots,2p}\mathrm{Perm}\left[\frac{1}{(1+w_{j}\bar{w}_{l})^{2}}\right]_{\begin{subarray}{c}j\neq k_{1},\ldots,k_{2p}\\ l\neq k^{\prime}_{1},\ldots,k^{\prime}_{2p}\end{subarray}}}\,.$ (S155)
In the end, we present the results for the one- and two-point correlators as examples.
For the one-point correlator, we have
${}_{\mathrm{NS}}\langle 0|\varepsilon(w,\bar{w})|\mathcal{C}_{\pm}\rangle$ $=\pm\sqrt{\frac{2-\sqrt{2}}{2}}\frac{1}{1+w\bar{w}}\,$, (S156)
@@
According to the fusion rule $\sigma\times\sigma\sim\mathbbm{1}+\varepsilon$, $(2n+1)$-point crosscap correlators of the $\sigma$ field vanish. For $2n$-point crosscap correlators, the calculation follows a similar logic to that of the $\varepsilon$ field. We express the full crosscap correlator as a combination of two Ishibashi crosscap correlators:
${}_{\mathrm{NS}}\langle 0|\prod_{j=1}^{2n}\sigma(w_{j},\bar{w}_{j})|\mathcal{C}_{\pm}\rangle=\sqrt{\frac{2+\sqrt{2}}{2}}{}_{\mathrm{NS}}\langle 0|\prod_{j=1}^{2n}\sigma(w_{j},\bar{w}_{j})|\mathbbm{1}\rangle\rangle_{\mathcal{C}}\pm\sqrt{\frac{2-\sqrt{2}}{2}}{}_{\mathrm{NS}}\langle 0|\prod_{j=1}^{2n}\sigma(w_{j},\bar{w}_{j})|\varepsilon\rangle\rangle_{\mathcal{C}}\,.$ (S159)
@@
${}_{\mathrm{NS}}\langle 0|\prod_{j=1}^{2n}\sigma(w_{j},\bar{w}_{j})|\mathbbm{1}\rangle\rangle_{\mathcal{C}}$ $=\sqrt{2^{n-1}\langle 0,0|\prod_{j=1}^{2n}\cos\frac{\vartheta}{2}(w_{j},\bar{w}_{j})\left(|\mathcal{O}_{+}\rangle+|\mathcal{O}_{-}\rangle\right)}\,$,
@@
$=\langle 0,-\frac{1}{2}\sum_{j=1}^{2n}\epsilon_{j}|\mathcal{O}^{0}_{\pm}\rangle\,\exp\left[-\frac{1}{4}\sum_{j=1}^{2n}\epsilon_{j}\,\sum_{l=1}^{2n}\epsilon_{l}\ln(w_{l}\bar{w}_{l})\right]\,$, (S165)
@@
$\phantom{=}\;\times\prod_{j=1}^{2n}|w_{j}|^{-2m\epsilon_{j}}\prod_{1\leq j<l\leq 2n}|w_{j}-w_{l}|^{\epsilon_{j}\epsilon_{l}/2}\prod_{1\leq j,l\leq 2n}\left(1+\frac{1}{w_{j}\bar{w}_{l}}\right)^{-\epsilon_{j}\epsilon_{l}/4}$